OpenStudy (anonymous):

Integrate of \[\frac{1+e ^{2x}}{e ^{x}}dx\]???

OpenStudy (anonymous):

just notice \[\frac{1+e^{2x}}{e^x}=\frac{1}{e^x}+\frac{e^{2x}}{e^x}=e^{x}+e^{-x}\]

OpenStudy (anonymous):

ok i got it. :) thanks!

OpenStudy (anonymous):

yw